OpenStudy (anonymous):

Given F(x) shown below, complete the equation for the inverse of F(x). If necessary, use the slash mark (/) for the division symbol. F(x)=x-1 f^-1(y)=_______

OpenStudy (anonymous):

y/1?

jimthompson5910 (jim_thompson5910):

f(x) = x-1 y = x-1 ... replace f(x) with y x = y -1 ... swap x and y now solve for y

jimthompson5910 (jim_thompson5910):

oh nvm about the swap, they just want you to solve y = x - 1 for x

OpenStudy (anonymous):

okay so its not y/1 its x/1?

OpenStudy (debbieg):

Right... weird notation, we saw this on another problem earlier.

OpenStudy (debbieg):

You have y=x-1 What do you do to solve that for x?

OpenStudy (anonymous):

you swap the numbers?

OpenStudy (debbieg):

I'm not sure what you mean by that. could you be more specific? Or just show the result.

OpenStudy (debbieg):

You either add, subtract, multiply or divide, {something} from both sides.

OpenStudy (anonymous):

ok you add the -1 and you get y/2

OpenStudy (debbieg):

hmmmm..... if you have y=x-1 How does adding 1 (which is what you mean: add 1, not add (-1)) to both sides get you y/2 anywhere?

OpenStudy (debbieg):

y+??=x-1+??

OpenStudy (anonymous):

y/0

OpenStudy (debbieg):

well, y/0 would be undefined... but more importantly, I'm puzzled as to how/why you are getting that?

OpenStudy (anonymous):

y=x-1 x=y-1 y=x+1

OpenStudy (debbieg):

\(y=x-1\) \(y+1=x-1+1\) add one to both sides \(y+1=x\) simplify on right side, since -1 +1=0 There is no division involved, but you keep wanting to divide y by stuff. Can you tell me why? I'm just trying to get at where your confusion is coming from.

OpenStudy (debbieg):

Well, yes.... that's how I would expect you to find the inverse. However, your teacher appears to want you to NOT swap x & y, but instead just solve the original equation for x. Which is a bit goofy, frankly. I would have done it the way YOU show above, getting y=x+1 as the inverse for y=x-1 That's really a notational matter though.

OpenStudy (anonymous):

so my final answer is...

OpenStudy (anonymous):

final answer is >.